OpenStudy (anonymous):

What is the y-intercept of the graph of 3x – 5y = 15 ? (0, –5) (0, 5) (0, –3) (0, 3)

OpenStudy (anonymous):

(0,-3)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

that's it?? ... you don't need to know how it goes ??!!

OpenStudy (anonymous):

Ummm , No thats fine

OpenStudy (ash2326):

Standard Equation in slope intercept form \[\frac{x}{a}+\frac{y}{b}=1\] a= x intercept b= y intercept We have \[3x-5y=15\] divide both sides by 15 we get \[\frac{x}{5}-\frac{y}{3}=1\] or \[\frac{x}{5}+\frac{y}{-3}=1\] so y intercept is -3